OpenStudy (anonymous):

Determine the exact value of sec (7 pi/6) @Mathematics

OpenStudy (anonymous):

sec theta = 1/cos theta, = 1/cos(7pi/6) <=> 1/-cos(1pie/6) = 1/-sqrt3/2 = 2/-sqrt3 <=> -2/sqrt3

OpenStudy (anonymous):

- 1/sqrt3